The electric potential at a point (x,y,z) is given by V = a + bx^n volt (a and b are positive constants).  The magnitude of electric field at (1,2,3) is 2 N / C and at (2,3,4) is 8 N / C. The electric field at (3,4,5) is? Pls answer with a brief explanation....
 

Grade:12th Pass

1 Answers

Samyak Jain
4 years ago
We know that electric field intensity at a point (x,y,z) is given by E = – [(\deltaV/\deltax)i + (\deltaV/\deltay)j + (\deltaV/\deltaz)k].
Hence, magnitude of electric field at any point is given by |dV/dr| in r direction.
Here, electric potential varies in x direction only, so we shall substitute dr as dx. Otherwise we need to put dr
as \deltax, \deltay, \deltaz separately to get electric field vector.
Now, V = a + b xn  \Rightarrow  dV/dx  =  bn xn – 1.
|dV/dx| at (1,2,3) is bn(1)n – 1  =  2 N/C  \Rightarrow  bn = 2         ...(1)
|dV/dx| at (2,3,4) is bn(2)n – 1  =  8 N/C                          ...(2)
Dividing (2) by (1), we get  (2)n – 1 = 4 = 22
\therefore n – 1 = 2  or  n = 3. Substitute the value of n in (1) to get b = 2/3.
We get V = a + (2/3) x3
So, |dV/dx| at (x,y,z) is 2 x2.
Thus, magnitude of electric field at (3,4,5) is   2.(3)2  =  18.